Literatur

Brannan D.A., Esplen M., Gray J.J.; Geometry, Cambridge University Press (1999) Coxeter H.S.M.; Unvergängliche Geometrie, Birkhäuser (1963) Jennings G.A.; Modern Geometry with applications, Springer (1994) Knörrer H.; Geometrie, Vieweg (1996) Koecher M.; Lineare Algebra und analytische Geometrie, Springer (1983) Neumann P.M., Stoy G.A., Thompson E.C.; Groups and Geometry, Oxford University Press (1994) Scriba C.J., Schreiber P.; 5000 Jahre Geometrie, Springer (2001)